OpenStudy (anonymous):

What real numbers are excluded from the domain of the following function? g(x)=7x+12/x^2+3x+7

OpenStudy (anonymous):

\[g(x)-\frac{ 7x+12 }{ x^2+3x+7}\]

OpenStudy (anonymous):

anything x which makes the denominator zero will be excluded. you have to solve:\[x^2+3x+7=0\]to find them.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

this one won't factor so use the quadratic formula to get the zeros.

OpenStudy (anonymous):

i am right now :)

OpenStudy (anonymous):

hint...you're going to run into trouble....

OpenStudy (anonymous):

....

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

can't figure it out! -_- help please?

OpenStudy (anonymous):

ok. the discriminant (the part of the quadratic under the square root sign) will be a negative number. Since you can't take the square root of a negative number there are no real solutions that make the denominator zero. i.e, there are no real numbers, x, which make the denominator zero.

OpenStudy (anonymous):

so, no matter what real values x takes, this function will accept it :) The domain is:\[x \in \mathbb{R} \]no real values are excluded.

OpenStudy (anonymous):

ok, thanks a lot!

OpenStudy (anonymous):

no problem